What is gravitational potential energy of the 135 kg object suspended 25 m above the earths surface?


1
Expert's answer
2021-01-29T20:01:26-0500

By the definition of the gravitational potential energy, we have:


"U=mgh=135\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot 25\ m=33075\ J."